If it takes Tasha 35 minutes to run 3.5miles on a steady speed, how many hours will it take her to run 25miles? Please provide workings. Thanks.
Answer by stanbon(75887)

If it takes Tasha 35 minutes to run 3.5miles on a steady speed, how many hours will it take her to run 25miles?
-------
In minutes::
m/25 miles = 35 min/3.5 miles
------
m = 25*10 = 250 min = (250/60) hrs = 25/6 hrs = 4 1/6 hrs = 4 hr 10 min
--------------
Cheers,
Stan H.
---------------